4 (25 pts) The vapor pressure of a solution containing 73.6 g glycerin (csHao,, Molar mass 92.094) in 120.7 g of ethanol (C H,OH, Molar mass 46.068) is 169.3 torr at 50.0°C. Calculate the vapor pressure of pure ethanol 40 0 Cassuming that glycerin is a nonvolatile, nonelectrolyte solute in ethanol. Explain2
Answer #1

vapor pressure of solution = moles fraction of solvent x vapor pressure of pure solvent

moles of ethanol = 120.7/46.068 = 2.62 mols

moles of glycerin = 73.6/92.094 = 0.80 mols

Total moles = 3.42 mols

moles fraction of ethanol = 2.62/3.42 = 0.766

vapor pressure of pure ethanol = 169.3/0.766 = 221.0 torr